Find the equation of a plane that is perpendicular to the vector 2i −2j +3k and passing through the point (3,4,−2)

The plane is given by:
Answer #1

\\*\text{The equation of a plane perpendicular to the vector }\vec n=a \hat i+b\hat j+c \hat k \\*\text{and passing through the point }(x_1,y_1,z_1)\text{ is:} \\*\text{\quad\quad\quad\quad\quad\quad} a ( x - x_1) + b (y - y_1) + c (z -z_1) = 0.

.

\\*\text{Here,} \\*\vec n=2\hat i-2\hat j+3\hat k\implies a=2,\quad b=-2,\quad c=3 \\*(x_1,y_1,z_1)=(3,4,-2)\implies x_1=3,\quad y_1=4,\quad z_1=-2

.

.

\\*\text{Hence, the required plane is} \\*2 ( x - 3) -2 (y -4) + 3 (z +2) = 0 \\*\implies 2x - 6 -2 y +8 + 3 z +6 = 0 \\*\implies \boxed{2x -2 y + 3 z +8= 0}